# Clustering Document Parts: Detecting and Characterizing Influence Campaigns From Documents

### v1

@@ Abstract
We propose a novel clustering pipeline to detect and characterize influence campaigns from documents. This approach clusters parts of document, detects clusters that likely reflect an influence campaign, and then identifies documents linked to an influence campaign via their association with the high-influence clusters. Our approach outperforms both the direct document-level classification and the direct document-level clustering approach in predicting if a document is part of an influence campaign. We propose various novel techniques to enhance our pipeline, including using an existing event factuality prediction system to obtain document parts, and aggregating multiple clustering experiments to improve the performance of both cluster and document classification. Classifying documents on the top of clustering not only accurately extracts the parts of the documents that are relevant to influence campaigns, but also capture influence campaigns as a coordinated and holistic phenomenon. Our approach makes possible more fine-grained and interpretable characterizations of influence campaigns from documents.
@@
Twitter
…Putin cleans up the bioweapons labs installed by the deep state… (44 tks)
…RT @EmmanuelMacron France strongly condemns Russia’s decision to wage war on Ukraine… (19 tks)
Forum
…a secret NATO laboratory for biological weapons…Biological weapons tests were carried out in the laboratories of this facility… (638 tks)
…[NATO] has blocked Ukraine’s plan to enter…Item 3: Ukraine was a pawn that the Westerners deliberately sacrificed to strengthen NATO… (703 tks)
News
…a NATO secret biological laboratory with biological weapons…The biological laboratory under the Azovstal plant in Marioupol in the so-called PIT-404 facility was built…In the laboratories of the facility, tests were carried out to create biological weapons… (1497 tks)
…Russia’s demand for neutrality…But NATO members said that Ukraine’s membership was at best a distant option… [The leader of the Ukrainian separatist region of Lugansk said he could hold a referendum on integration into Russia,] a decision immediately criticized by Kiev…(1152 tks)
@@ 1 Introduction
An influence campaign is a coordinated and strategic effort to shape and manipulate the perceptions of a target audience about certain things or issues over a period of time Martin et al. (2023). It can be organized by an individual, organization, or government for various purposes, such as promoting a specific public image, product, policy, or political narrative, and through various channels, including traditional media and online platforms. Consequently, detecting an influence campaign requires holistic evaluations and the use of multiple indicators, such as the social network (des Mesnards and Zaman, 2018), that point to a collective effort with a shared motive that aims to impact public opinions in a certain way. Accurate and reliable detection typically involves extensive manual verification by domain experts (Martin et al., 2023).
@@
In this paper, we propose a novel clustering-based pipeline to help detect and characterize influence campaigns from documents. Unlike the typical document-level classification approach discussed above, the pipeline predicts influence campaigns directly on the cluster level, i.e., if a cluster of document parts present an influence campaign (a high-influence cluster). From there, the pipeline further predicts whether any document associated with a high-influence cluster is part of the influence campaign via a dynamic projection procedure. As a result, our pipeline is capable of handling the two aspects of the influence campaign detection task: capturing influence campaigns as a holistic phenomenon, and predicting documents that are part of an influence campaign. Since influence campaigns are captured by clusters of document parts and the documents predicted to engage in an influence campaign are projected from these clusters, our pipeline enables fine-grained and interpretable characterizations of influence campaigns from documents. The specific contributions of this paper are as follows.
@@
This paper is structured as follows. We review related works in Sec 2 and motivate and explain the algorithmic idea underlying our novel clustering pipeline in Sec 3. We describe the influence campaign detection dataset we test our pipeline on in Sec 4 and the experiments in Sec 5. The results are discussed in Sec 6. The paper concludes in Sec 7.
@@ 2 Related Work
There have been very few studies in the existing literature that approach influence campaigns in the general sense as we define it in Sec 1. The influence campaigns studied in most previous research (des Mesnards and Zaman, 2018; Luceri et al., 2023; Martin et al., 2023) are political influence campaigns, or some closely related political influence operations that may be an influence campaign, such as the spreading of mis/dis-information (Ferrara, 2017; Álvaro Figueira and Oliveira, 2017; Rubin, 2017; Addawood et al., 2019; Barrón-Cedeño et al., 2019; Nogara et al., 2022; Sakketou et al., 2022; Malik MSI, 2023).
The most common detection method relevant for influence campaigns is bot detection in social networks (Davis et al., 2016; Badawy et al., 2018; des Mesnards and Zaman, 2018; Himelein-Wachowiak et al., 2021; Hajli et al., 2022; Rossetti and Zaman, 2022). For text-based influence campaign detection, various NLP methods have been explored. For example, a recent study leverages LLMs (Luceri et al., 2023) to predict if a tweet is part of an known influence campaign. Other studies relevant for influence campaigns utilize various sources of linguistic features (e.g., lexicon counts, ngrams, word embedding) to train or fine-tune different models (e.g., BERT, graph neural network, decision tree) with a goal to detect propagandistic, deceptive, or misleading information (Addawood et al., 2019; Barrón-Cedeño et al., 2019; Sakketou et al., 2022; Malik MSI, 2023). To the best of our knowledge, we are not aware of any study that aims to detect influence campaigns on the cluster level.
Several twitter datasets have been used by recent studies on detecting influence campaigns (des Mesnards and Zaman, 2018; Luceri et al., 2023), such as the 2016 US election dataset (Littman et al., 2016), data from Twitter’s Information Operations archive [^2] E.g., https://blog.twitter.com/en_us/topics/company/2020/2020-election-changes, and Russian troll accounts for 2016 US election released by the U.S. Congress (Addawood et al., 2019). There are also other relevant datasets in other media forms, such as FACTOID (Sakketou et al., 2022) collected from Reddit and Proppy Barrón-Cedeño et al. (2019) collected from news articles. We note that for all of these datasets that come with labels, the labels are typically created on the basis of some simplistic association. For example, tweets are assumed to be linked to an influence campaign if they come from Russian troll accounts. We do not know of any publicly available influence campaign datasets that contain more than one media types.
@@ 3 Pipeline: The Algorithmic Idea
Given the coordinated nature of influence campaigns, an influence campaign can be thought of as a cluster of documents that spread a certain theme. Our pipeline follows exactly this intuition and transforms the task of influence campaign detection into one that detects clusters that are highly likely to reflect an influence campaign (high-influence clusters). Then the next step naturally becomes how to accurately select documents (high-influence documents) associated with the high-influence clusters that reflect an influence campaign, assuming the clusters may contain some noise or false positives. More concretely, our pipeline consists of the following four steps.
@@
In a pre-processing step, we start out by extracting parts from a document. In this paper, we experiment with the spans that represent what the author expresses certain belief in; sentences; and the whole document as a default.
@@
Given a set of documents, the pipeline clusters the document parts. Clustering parts of documents not only creates a complex connection network among documents via their semantically related parts, but also presents an effective workaround for long document information retrieval using unsupervised clustering algorithms (Mekontchou et al., 2023).
@@
During the training time, the pipeline takes as inputs a set of documents, each of which is annotated with a binary label: the document is or is not part of an influence campaign. The concept of high-influence clusters is defined by a user-given value $\alpha$, denoting the percentage of document parts in a cluster linked to documents from an influence campaign.
Intuitively, $\alpha$ should be far greater than 0.5 to align with the heuristic that a high-influence cluster should be dominated by document parts from documents that are part of an influence campaign. The assumption here is that parts of documents with a link to an influence campaign are unlikely to be clustered together unless they are related to some aspect or surface theme of the influence campaign. Since high-influence clusters may be rare for a given clustering experiment when the majority of documents are “innocent”, we propose doing multiple clustering experiments and aggregating the resulting clusters together as a way to generate more data to train a classifier for high-influence clusters, a novel data augmentation technique for cluster-level classification. In the paper, we set $\alpha=0.95$ for reasons to be explained later.
At inference time, the pipeline deploys the pre-trained classifier to detect high-influence clusters.
@@
High-influence documents are ones with connections to high-influence clusters, meaning their parts occur in at least one high-influence clusters. Formally, we set a threshold $\beta$, a ratio for the number of high-influence clusters, which denotes the minimum number of times the parts of a document must occur in any high-influence clusters to qualify the document as a high-influence document. The threshold $\beta$ is a result of the threshold $\alpha$ set in the previous step. Suppose (at training time) we project the defined high-influence clusters to high-influence documents and $\alpha=1$, the precision of identifying documents related to an influence campaign is guaranteed to be 100%, because no false positives exist in the high-influence clusters by definition. It turns out in practice that having an error term $1-\alpha$ ($\alpha\neq 1$) in the definition of high-influence clusters facilitates discovery of more high-influence clusters and documents, i.e., a great improvement in recall at a small cost of precision, ultimately leading to a better F1. We wish to come up with a module in our current system that predicts an optimal threshold $\beta$ in the future.
In summary, we have introduced the notion of high-influence clusters based on parts of document and the notion of high-influence documents based on their association with high-influence clusters. As a response, we propose two thresholds ($\alpha$ and $\beta$) to regulate the number of false high-influence documents our system may end up selecting from high-influence clusters. We also propose the aggregation of clustering experiments, instead of hard fine-tuning for an optimal clustering experiment, to enhance model performance. In what follows, we show that our approach can easily and significantly outperform direct document-level classification, in an apples-to-apples comparison, when it comes to detecting influence campaigns from documents.
@@ 4 Data
We use data collected during a large research program with a DARPA INCAS project [^3] https://www.darpa.mil/program/influence-campaign-awareness-and-sensemaking. We expect the data to be made public after the end of the research program. We use this dataset as we are not aware of any other datasets that have expert-verified annotations indicating if a collection of documents contain influence campaigns.
The data contains four piles of online posts published during January 31 to June 30, 2022. Each pile is a collection of documents in six media forms, namely, Twitter, Forum, News, Blog, Reddit, and Other. Two of the four piles contain documents that engage in an influence campaign that spreads disinformation related to Ukraine bioweapons conspiracy theory (see Table 1), whereas the other two contain Ukraine-related documents with no links to any known influence campaigns. Lexical search was used to facilitate the collection of the data. Over 99% of the documents that participate in the bioweapons influence campaign use words like “biolab” and “biological weapons”, but slightly less than 3% of the documents unrelated to the campaign mention these terms.
The majority of these documents are written in French, typically accompanied with an English translation, and the rest are in English. We choose to work on the translated French portion of the data. This portion has over 8 times more documents than the English counterpart, but with a significantly smaller portion of documents linked to an influence campaign, less than 8%. We believe this represents a more realistic setting for detecting influence campaigns from documents.
We randomly split the translated French data into two parts, the train and test sets. We split at the document level, with a ratio of 80/20, as shown in Table 2. Appendix A provides further details about the distribution of media forms and average document length in the data.
@@
[…]
@@ 5.1 Task description
As demonstrated in Sec 1, the real challenge for detecting influence campaigns is how to capture the phenomenon of influence campaigns. Making the detection of influence campaigns a binary classification task, i.e., to predict whether a text or a document is part of an influence campaign, is not only less realistic but also probably doomed; such a classification approach cannot survive the constantly shifting and evolving nature of influence campaigns as a social phenomenon.
Nevertheless, to comprehensively evaluate our new pipeline requires a large-scale dataset annotated on the document collection level, indicating if a collection of documents present an influence campaign. Since the dataset described in Sec 4 is the only dataset in this regard, this comprehensive evaluation cannot be possible.
@@ 5.2 Baselines
We train two direct document-level classifiers (Direct-document), using fully connected feedforward neural networks (FNN) and XGBoost algorithm (Chen and Guestrin, 2016), as the baselines to compare with our approach. XGBoost is an optimized gradient boosting (Friedman, 2001) system using tree ensembles that achieve state-of-the-art results on many real-world machine learning challenges. We refrain from using any word-embedding-based machine learning models to prevent models from learning from general lexical data, which precludes the use of models such as LLMs, BERT, LSTM, and so on. We use 95 general linguistic features, extracted by an open-sourced corpus linguistics tool (Wang, 2021), to train the models. These features are mostly based on the work of Biber (1988, 2006) and have been developed over decades to suit general text analysis. We also add “number of words” as a feature to factor in document length, particularly for short documents (say, a tweet), which may not see occurrence of many features at all due to length limitation. The extracted features are mostly normalized frequency counts. More details about the model parameters and these features are given in Appendix B.
In addition, we apply our pipeline on the document level (Document-level), namely, clustering the entire documents, as additional baselines to emphasize the importance of clustering documents parts. The experimental setup for document-level clustering is identical to the setup for our approach.

### v2

@@ Abstract
We propose a novel clustering pipeline to detect and characterize influence campaigns from documents. This approach clusters parts of document, detects clusters that likely reflect an influence campaign, and then identifies documents linked to an influence campaign via their association with the high-influence clusters. Our approach outperforms both the direct document-level classification and the direct document-level clustering approach in predicting if a document is part of an influence campaign. We propose various novel techniques to enhance our pipeline, including using an existing event factuality prediction system to obtain document parts, and aggregating multiple clustering experiments to improve the performance of both cluster and document classification. Classifying documents after clustering not only accurately extracts the parts of the documents that are relevant to influence campaigns, but also captures influence campaigns as a coordinated and holistic phenomenon. Our approach makes possible more fine-grained and interpretable characterizations of influence campaigns from documents.
@@
Twitter …Putin cleans up the bioweapons labs installed by the deep state… (44 tks) …RT @EmmanuelMacron France strongly condemns Russia’s decision to wage war on Ukraine… (19 tks)
Forum …a secret NATO laboratory for biological weapons…Biological weapons tests were carried out in the laboratories of this facility… (638 tks) …[NATO] has blocked Ukraine’s plan to enter…Item 3: Ukraine was a pawn that the Westerners deliberately sacrificed to strengthen NATO… (703 tks)
News …a NATO secret biological laboratory with biological weapons…The biological laboratory under the Azovstal plant in Marioupol in the so-called PIT-404 facility was built…In the laboratories of the facility, tests were carried out to create biological weapons… (1497 tks) …Russia’s demand for neutrality…But NATO members said that Ukraine’s membership was at best a distant option… [The leader of the Ukrainian separatist region of Lugansk said he could hold a referendum on integration into Russia,] a decision immediately criticized by Kiev…(1152 tks)
@@ 1 Introduction
Inspired by Martin et al. (2023) and Luceri et al. (2023), we define an influence campaign as a coordinated and strategic effort to shape and manipulate the perceptions of a target audience about certain things or issues over a period of time. It can be organized by an individual, organization, or government for various purposes, such as promoting a specific public image, product, policy, or political narrative. It can be carries out through various channels, including traditional media and online platforms. Consequently, detecting an influence campaign requires holistic evaluations and the use of multiple indicators, such as the social network (des Mesnards and Zaman, 2018), that point to a collective effort with a shared motive that aims to impact public opinions in a certain way. Accurate and reliable detection typically involves extensive manual verification by domain experts, taking into account both textual and non-textual information (Martin et al., 2023).
@@
In this paper, we propose a novel text-only clustering-based pipeline to help detect and characterize influence campaigns from documents. Unlike the typical document-level classification approach discussed above, the pipeline predicts influence campaigns directly on the cluster level, i.e., it predicts whether a cluster of document parts present an influence campaign (a high-influence cluster). From there, the pipeline further predicts whether any document associated with a high-influence cluster is part of the influence campaign via a dynamic projection procedure. As a result, our pipeline is capable of handling the two aspects of the influence campaign detection task: capturing influence campaigns as a holistic phenomenon, and predicting documents that are part of an influence campaign. Since influence campaigns are captured by clusters of document parts and the documents predicted to engage in an influence campaign are projected from these clusters, our pipeline enables fine-grained and interpretable characterizations of influence campaigns from documents. The specific contributions of this paper are as follows.
@@
This paper is structured as follows. We review related works in Sec 2 and motivate and explain the algorithmic idea underlying our novel clustering pipeline in Sec 3. We describe the influence campaign detection dataset we test our pipeline on in Sec 4 and the experiments in Sec 5. The results are discussed in Sec 6. The paper concludes in Sec 7. We release our code at https://github.com/jaaack-wang/detect-influence-campaigns.
@@ 2 Related Work
There have been very few studies in the existing literature that approach influence campaigns in the general sense as we define it in Sec 1. The influence campaigns studied in most previous research (des Mesnards and Zaman, 2018; Luceri et al., 2023; Martin et al., 2023) are political influence campaigns, or some closely related political influence operations that may be an influence campaign, such as the spreading of mis/dis-information (Ferrara, 2017; Álvaro Figueira and Oliveira, 2017; Rubin, 2017; Addawood et al., 2019; Barrón-Cedeño et al., 2019; Nogara et al., 2022; Sakketou et al., 2022; Malik et al., 2023).
The most common detection method relevant for influence campaigns is the detection of bots in social networks (Davis et al., 2016; Badawy et al., 2018; des Mesnards and Zaman, 2018; Himelein-Wachowiak et al., 2021; Hajli et al., 2022; Rossetti and Zaman, 2022). For text-based influence campaign detection, various NLP methods have been explored. For example, a recent study leverages LLMs (Luceri et al., 2023) to predict if a tweet is part of an known influence campaign. Other studies relevant for influence campaigns utilize various sources of linguistic features (e.g., lexicon counts, ngrams, word embedding) to train or fine-tune different models (e.g., BERT, graph neural network, decision tree) with a goal to detect propagandistic, deceptive, or misleading information (Addawood et al., 2019; Barrón-Cedeño et al., 2019; Sakketou et al., 2022; Malik et al., 2023). To the best of our knowledge, we are not aware of any study that aims to detect influence campaigns on the cluster level.
Several twitter datasets have been used by recent studies on detecting influence campaigns (des Mesnards and Zaman, 2018; Luceri et al., 2023), such as the 2016 US election dataset (Littman et al., 2016), data from Twitter’s Information Operations archive [^2] E.g., https://blog.twitter.com/en_us/topics/company/2020/2020-election-changes, and Russian troll accounts for 2016 US election released by the U.S. Congress (Addawood et al., 2019). There are also other relevant datasets in other media forms, such as FACTOID (Sakketou et al., 2022) collected from Reddit and Proppy Barrón-Cedeño et al. (2019) collected from news articles. We note that for all of these datasets that come with labels, the labels are typically created on the basis of some simplistic association or assumption. For example, tweets are assumed to be linked to an influence campaign if they come from Russian troll accounts (Luceri et al., 2023). To the best of our knowledge, we do not know of any publicly available influence campaign datasets that contain more than one media type.
@@ 3 Pipeline: The Algorithmic Idea
Given the coordinated nature of influence campaigns, an influence campaign can be thought of as a cluster of documents that spread a certain theme aimed to influence the target audience. Our pipeline follows exactly this intuition and transforms the task of influence campaign detection into one that detects clusters that are highly likely to reflect an influence campaign (i.e., high-influence clusters). Then the next step naturally becomes how to accurately select documents (i.e., high-influence documents) associated with the high-influence clusters that reflect an influence campaign, assuming the clusters may contain some noise or false positives.
More concretely, our pipeline consists of the following four steps.
@@
In a pre-processing step, we start out by extracting parts from a document. In this paper, we experiment with three types of document parts: the multi-word text spans that represent what the author expresses certain belief in (see Sec 5.3); sentences; and the whole document.
@@
Given a set of documents, the pipeline clusters the document parts. Clustering parts of documents not only creates a complex connection network among documents via their semantically related parts, but also presents a general and effective workaround for long document information retrieval using unsupervised clustering algorithms (Mekontchou et al., 2023).
@@
At training time, the pipeline takes as inputs a collection of documents, each of which is annotated with a binary label: the document is or is not part of an influence campaign. The fact that our pipeline requires annotated documents during training highlights that it is a supervised approach. The concept of high-influence clusters is defined by a user-given threshold $\alpha$, denoting the minimum percentage of document parts in a cluster linked to documents from an influence campaign for the cluser to count as a high-influence cluster.
Intuitively, $\alpha$ should be set far greater than 0.5 to align with the heuristic that a high-influence cluster should be dominated by document parts from documents that are part of an influence campaign. The assumption here is that parts of documents with a link to an influence campaign are unlikely to be clustered together unless they are related to some aspect or surface theme of the influence campaign. Since high-influence clusters may be rare or even absent for a given clustering experiment for which the majority of documents are “innocent” and do not reflect an influence campaign, we propose doing multiple clustering experiments and aggregating the resulting clusters together as a way to generate more data to train a classifier for high-influence clusters. This approach is a novel data augmentation technique for cluster-level classification. In this paper, we set $\alpha=0.95$. We set $\alpha\neq 1$ as a trade of the precision and recall for discovering high-influence clusters, since allowing a small error term $1-\alpha$ ($\alpha\neq 1$) in the definition of high-influence clusters facilitates discovery of more high-influence clusters, i.e., a great improvement in recall at a small cost of precision, ultimately leading to a better F1. Note that $\alpha$ is only set and used at the training time.
At inference time, the pipeline deploys the pre-trained classifier to detect high-influence clusters by predicting the likelihood of a cluster being a high-influence cluster.
@@
High-influence documents are documents with connections to high-influence clusters, meaning at least some of their parts occur in at least one high-influence clusters. Formally, we set a threshold $\beta$, a ratio for the number of high-influence clusters, which denotes the minimum number of times parts of a document that must occur in any high-influence clusters to qualify the document as a high-influence document. The threshold $\beta$ is used to control the number of false positives (i.e., documents with no link to an influence campaign occurring in high-influence clusters) introduced by the threshold $\alpha$ set in the previous step. We wish to come up with a module in our current system that predicts an optimal threshold $\beta$ in the future.
In summary, we have introduced the notion of “high-influence clusters” based on parts of documents and the notion of “high-influence documents” based on their association with high-influence clusters. We propose two thresholds ($\alpha$ and $\beta$) to regulate the number of false-positive high-influence documents our system may end up selecting from high-influence clusters. The threshold $\alpha$ is used for training a high-influence cluster classifier, whereas the threshold $\beta$ is used at the time of classifying high-influence documents. We also propose the aggregation of clustering experiments, instead of hard fine-tuning for an optimal clustering experiment, to reliably enhance model performance.
In what follows, we show that our approach can easily and significantly outperform direct document-level classification, in an apples-to-apples comparison, when it comes to detecting influence campaigns from documents.
@@ 4 Data
We use data collected during a large research program, DARPA INCAS project [^3] https://www.darpa.mil/program/influence-campaign-awareness-and-sensemaking. We expect the data to be made public after the end of the research program. We use this dataset as we are not aware of any other datasets that have expert-verified annotations indicating if a collection of documents contain influence campaigns.
The data contains four piles of online posts published during January 31 to June 30, 2022. Each pile is a collection of documents in six media forms, namely, Twitter, Forum, News, Blog, Reddit, and Other. Two of the four piles contain documents that engage in an influence campaign that spreads disinformation related to Ukraine bioweapons conspiracy theory (see Table 1), whereas the other two contain Ukraine-related documents with no links to any known influence campaigns. Lexical search was used to facilitate the collection of the data. Over 99% of the documents that participate in the bioweapons influence campaign use words like “biolab” and “biological weapons”, but slightly less than 3% of the documents unrelated to the campaign mention these terms. That means that any content-based text classifier, whether rule-based or neural, will overfit this dataset by capturing the related keywords. We avoid training such classifiers, as (1) we are more interested in developing a potentially general approach that can both detect and characterize influence campaigns from documents; and (2) content-based text classification for the influence campaign detection task arguably cannot be a general approach nor can it make predictions at beyond the document level to capture the phenomenon of influence campaign.
The majority of these documents are written in French, typically accompanied by an English translation, and the rest are in English. We choose to work on the translated French portion of the data. This portion has over 8 times more documents than the English subcorpus, but with a significantly smaller portion of documents linked to an influence campaign, less than 8%. We believe this represents a more realistic and challenging setting for detecting influence campaigns from documents.
Given the overall small size of the dataset, we randomly split the French data into two parts, the train and test sets, but the train set can be further split for training and validation where needed. We split at the document level, with a ratio of 80/20, as shown in Table 2. Appendix A provides further details about the distribution of media forms and average document length in the data.
@@
Table 2: Statistics of the train and test sets (disjoint). Numbers inside parentheses show the number and percentage of documents or document parts linked to an influence campaign. TargetsAT: targets the Author herself believes to be True. TargetsALL: all belief targets, regardless of the belief holder and the commitment level.
@@ 5.1 Task description
As argued in Sec 1, the real challenge for detecting influence campaigns is how to capture the phenomenon of influence campaigns. Making the detection of influence campaigns a binary classification task, i.e., to predict whether a text or a document is part of an influence campaign, is not only less realistic but also probably doomed; such a classification approach cannot survive the constantly shifting and evolving nature of influence campaigns as a dynamic social phenomenon.
Nevertheless, to comprehensively evaluate our new pipeline requires a large-scale dataset annotated on the document collection level, indicating if a collection of documents presents an influence campaign. Since the dataset described in Sec 4 is the only dataset in this regard, this comprehensive evaluation cannot be possible.
@@ 5.2 Baselines
We train two direct document-level classifiers (Direct-document), using fully connected feedforward neural networks (FNN) and XGBoost algorithm (Chen and Guestrin, 2016), as the baselines to compare with our approach. XGBoost is an optimized gradient boosting (Friedman, 2001) system using tree ensembles that achieve state-of-the-art results on many real-world machine learning challenges. We refrain from using any word-embedding-based or content-based machine learning models to prevent models from learning from general lexical data, which precludes the use of models such as LLMs, BERT, LSTM, and so on. We use 95 general linguistic features, extracted by an open-sourced corpus linguistics tool (Wang, 2021), to train the models. These features are mostly based on the work of Biber (1988, 2006) and have been developed over decades to suit general text analysis. We also add “number of words” as a feature to factor in document length, particularly for short documents (say, a tweet), which may not see the occurrence of many features at all due to length limitation. The extracted features are mostly normalized frequency counts. More details about the model parameters and these features are given in Appendix B.
In addition, we apply our pipeline on the document level (Document-level), namely, clustering the entire documents, as an additional baseline to emphasize the importance of clustering document parts. The experimental setup for document-level clustering is identical to the setup for our approach based on document parts.